OpenStudy (anonymous):

Help with writing expressions: Neil is selling candy bars at the ball game and on average, manages to sell one bar every 2 minutes.

OpenStudy (anonymous):

OK what's the question?

OpenStudy (anonymous):

By reading that I need to come up with an expression

OpenStudy (phi):

write it as a ratio

OpenStudy (anonymous):

OK... Well I guess you could say that x= time in minutes y= candy bars sold in total If he sells 1 bar every 2 minutes, that's the same as 1/2 candy bar every minute. So y=(1/2)x

OpenStudy (anonymous):

So what if the guy stated with 64

OpenStudy (anonymous):

started with 64??

OpenStudy (anonymous):

Yeah I got the expression is -1/2x+64

OpenStudy (phi):

I would write \[ \frac{ 1 \text{ candy bar }}{2 \text{ minutes }} \]

OpenStudy (phi):

now what do you want to do with this ?

OpenStudy (phi):

If he has sold 64 candy bars then you can set up a ratio: \[ \frac{ 1 \text{ candy bar }}{2 \text{ minutes }} = \frac{ 64 \text{ candy bar }}{x \text{ minutes }} \]